OpenStudy (anonymous):

solve x2+14x+49=0 (x+7(x+7)? I don't know if it's rite

jimthompson5910 (jim_thompson5910):

Good, you factored x^2+14x+49 into (x+7)(x+7), now use the zero product property to solve x^2+14x+49 = 0 (x+7)(x+7) x+7 = 0 or x+7 = 0 .... Use the zero product property here x = -7 or x = -7 So the only solution is x = -7

OpenStudy (anonymous):

Oh thanks!

jimthompson5910 (jim_thompson5910):

You're welcome